You are the chief financial officer for a firm that sells digital music players. Your firm has the

following average cost schedule;

Quantity Average total cost

500players $300


501 301

Your current level of production is 500 devices, all of which have been sold. Someone calls,

desperate to buy one of your music players. The caller offers you $450 for it. Should you accept

the offer? Why or why not?


1
Expert's answer
2021-04-26T19:34:52-0400

At the current level of production of 500 devices and the price of $450 the total profit is:

"TP = (450 - 300)\u00d7500 = 75,000."

At the level of 501 devices produced the total profit is:

"TP = (450 - 301)\u00d7501 = 74,649."

As average cost increases and total profit decreases as a result of one more unit produced, then you shouldn't accept the offer.